OpenStudy (anonymous):

y=-3cos πx what is the period ?

OpenStudy (anonymous):

the period is \[\frac{ 2\pi }{ \pi }=2\]

OpenStudy (anonymous):

period is \[\frac{ 2\pi }{ b }\] whatever that b might be

OpenStudy (anonymous):

answer key says that it's \[\frac{ 3\pi }{ 2 }\]

OpenStudy (anonymous):

are you sure?

OpenStudy (anonymous):

yes , that's why im confused too

OpenStudy (anonymous):

then discard the book answer :D

OpenStudy (anonymous):

ohh that's true too :P would the graph be correct?

OpenStudy (anonymous):

from the picture above yea i'd say so